OpenStudy (anonymous):

What's the solution for H^2 = -X^2 + 2x +1

OpenStudy (tylerd):

can you give any more info?

OpenStudy (anonymous):

squaring both sides of the equation and solving for H we have... 2x^2+H^2=x^2+2x+1 H^2=-x^2+2x+1 H=???

OpenStudy (tylerd):

you have one equation and 2 unknowns, im not sure how you could solve it.

OpenStudy (tylerd):

nvm that wouldnt work

OpenStudy (anonymous):

This math is hard as hell haha. I should have gone to college so I could be more help haha

OpenStudy (tylerd):

Ive never seen a problem like this and im finishing up the calculus series at the moment.

OpenStudy (anonymous):

I really do appreciate your help bro

OpenStudy (tylerd):

what topic is this? or chapter, or subject?

OpenStudy (anonymous):

3.4 pre cal

OpenStudy (anonymous):

Or math 126

OpenStudy (tylerd):

is this completing the square?

OpenStudy (anonymous):

Apparently it's real 0's of polynomials